TWiki home Point 1 Point 2 Point 3 Point 4 Point 5 Point 6 Point 7 Point 8 Minutes


Created by Main.BorisBellesia on 2007-07-30 - 19:29

Short Circuit Test in UJ16-RR17

PDF RAT_20070730.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint1x2007x07x30x19x27/RAT_20070730.pdf

Comments:

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-07-26 - 17:12

Short Circuit Test RR17-UJ16 (also UA23-UA27)

PDF RAT_2007070626.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint1x2007x07x26x17x10/RAT_2007070626.pdf

Comments:

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-07-24 - 18:20

Short Circuit Tests in RR17-UJ16 (also UA23-UA27)

PDF Circuit_groups.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint1x2007x07x24x18x17/Circuit_groups.pdf

PDF sct_RR17_v003.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint1x2007x07x24x18x17/sct_RR17_v003.pdf

PDF RAT_2007070624.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint1x2007x07x24x18x17/RAT_2007070624.pdf

Comments:

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-07-20 - 17:56

Short Circuit Test UJ16-RR17 (also UJ33 - UA27/UA23)

PDF RAT_2007070620.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint1x2007x07x20x17x51/RAT_2007070620.pdf

Heat run in UJ33

The 24hr heat run has been successfully performed. The re-routing of the cables was successful. The temperatures measured in the UJ33 are now in line with the ones recorded in the other LHC short circuit test zones and the so the temperature inside the FGC and 600A EE racks. The detailed summary of the heat run will be presented at the HCC meeting on the 2nd of August.

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-07-18 - 20:04

Short Circuit Test in RR17-UJ16 (also UJ33 + UA23/UA27)

PDF RAT_2007070618.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint1x2007x07x18x20x02/RAT_2007070618.pdf

PDF sct_UA23+RR17+UJ33_v001.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint1x2007x07x18x20x02/sct_UA23+RR17+UJ33_v001.pdf

Comments:

Edit, attach file or add comment


Created by Main.MirkoPojer on 2007-07-10 - 15:17

Warm Magnets Commissioning of Point 7 - Minutes from meeting_10/07/07

PDF RAT_20070710_WM7.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint7x2007x07x10x15x13/RAT_20070710_WM7.pdf

Comments:

Edit, attach file or add comment


Created by Main.MirkoPojer on 2007-07-10 - 09:31

Beam Dump System - Installation and Commissioning Meeting _ 03/07/07

PDF RAT_20070703.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPointx2007x07x10x09x30/RAT_20070703.pdf

Comments:

Edit, attach file or add comment


Created by Main.MariaPazCasasLino on 2007-07-06 - 12:25

Short Circuit Test UA23. Minutes 06-07-2007

PDF sct_UA23_v004.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint2x2007x07x06x12x23/sct_UA23_v004.pdf

PDF SOC_UA23-UJ23_06-07-2007.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint2x2007x07x06x12x23/SOC_UA23-UJ23_06-07-2007.pdf

PDF RAT_20070706.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint2x2007x07x06x12x23/RAT_20070706.pdf

Comments:

Edit, attach file or add comment


Created by Main.AntonioVergara on 2007-07-05 - 16:38

Powering Tests Sector 7-8 - Thursday 05.07.2007

Microsoft Excel Spreadsheet PlanningW27_050707.xls:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x07x05x16x36/PlanningW27_050707.xls

Comments:

Edit, attach file or add comment


Created by Main.MirkoPojer on 2007-07-05 - 00:13

Powering tests Sector 7-8 _ 04/07/2007

Morning shift activity.

  • RSD2.A78B2 done up to PCS4
  • RQS.A78B2 wrong polarity
  • RQS.R7B1 wrong polarity
  • RQTL10.R7B1 wrong polarity
  • RSD1.A78B1 high noise
  • RSD1.A78B2 card problems
  • RSF2.A78B1 done up to PCS2
  • RSF2.A78B2 tripped during PCS1
  • RSF1.A78B2 done up to PCS3

Performance test done on: RQS.L8B1 RCS.A78B2 RCS.A78B1

24h run stopped at 11:30.

Afternoon shift.

The following circuits have been testes up to PCS4:
  • RQS.A78B2
  • RQS.R7B1
  • RQTL10.R7B1
  • RSF1.A78B2
  • RSF2.A78B1
  • RSF2.A78B2

The following circuits are affected by a problem in a QPS card: RSF1.A78B1 RSD1.A78B1 RSD1.A78B2

RQ6.R7B1/B2 were tested but tripped due to QPS threshold. Reiner needs to increase the threshold there. In agreement with MPP, it has been decided to raise their threshold to 50 mV.

Activity for tomorrow, Thursday 5th.

  • condamnation of all coverters of point 7, to allow ElQA to start their tests there (RQ6, in DFBMH, is on the same electrical depart as other circuits of DFBAN);
  • Reiner will work on the QPS system for RSD1.A78B1/B2 and RSF1.A78B1 tomorrow morning around 9.00 (change of detection card required); as soon as he finishes, those circuits will be powered up to PCS4;
  • as soon as the three circuits above are over, condamnation of all converters of DFBAO;
  • as soon as Reiner finishes the intervention on RQ6 (probably early afternoon), the corresponding converters will be de-condamned and RQ6 powered for PCS4 (THIS HAS TO BE DONE IN ACCORDANCE WITH ElQA FOR WHAT MENTIONED ABOVE ON THE SAME ELECTRICAL DEPART!!!); the converters will be condamned again right after.

  • ElQA activity:
    • matching section to be completed;
    • RB to qualify (after disconnection by Philip Denis of leads in point 7)
    • RQF/D qualification
    • 60 A COC

    • DFBAN ready from the morning
    • DFBAO ready from the afteernoon

  • Installation of passive heating system on one 600A of DFBAN (Vincent Fontanive)

  • DFBXG.3L8 will be cold-tested tomorrow afternoon.

Edit, attach file or add comment


Created by Main.AntonioVergara on 2007-07-04 - 16:47

Powering Tests Sector 7-8 - Test Progress Wednesday 04.07.2007 Afternoon

Microsoft Excel Spreadsheet PlanningW27_040707_afternoon.xls:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x07x04x16x26/PlanningW27_040707_afternoon.xls

Comments:

Edit, attach file or add comment


Created by Main.MariaPazCasasLino on 2007-07-04 - 11:11

Short Circuit Test in UA23. Minutes 04-07-2007

PDF RAT_20070704.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint2x2007x07x04x10x58/RAT_20070704.pdf

PDF sct_UA23_v003.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint2x2007x07x04x10x58/sct_UA23_v003.pdf

PDF SOC_8h_run_(06-07-2007).pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint2x2007x07x04x10x58/SOC_8h_run_(06-07-2007).pdf

Comments:

Note that next Thursday 5th July a pre-test for the 8h heat run will be carried out. (Start at 14:00h)

Note that next Friday 6th July a 8h heat run will take place (Circuits < 13kA)

-- Main.MariaPazCasasLino - 04 Jul 2007

Edit, attach file or add comment


Created by Main.UnknownUser on 2007-07-03 - 21:13

Report from Tuesday 3rd July 2007 - Powering Tests Sector 78

Microsoft Excel Spreadsheet PlanningW27_040707.xls:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x07x03x19x54/PlanningW27_040707.xls

Tests carried out on Tuesday 03.07.2007

  • 600A Correctors
    • PCS completed OK for: RCS.A78B2, RCS.A78B1, RQTD.A78B1, RQTD.A78B2, RSD2.A78B1
    • PCS.3 (100A) completed OK for RSF1.A78B1
    • PLI1 (50A), PLI2 (100A) and PNO (550A) for RQS.L8B1 - Test OK
    • ElQA, PCC and PIC2 completed for RQS.A78B2, RQS.R7B1 and RQTL10.R7B1 - Test OK
    • Stability tests on RQT12.L8B1 completed - Test OK
    • Stability tests on RQTF.A78B1 ongoing

  • Manual Tracking tests for RQF.178, RQD.A78 and RB.A78 completed

  • LSA tracking tests for RQF.178, RQD.A78, RB.A78, RQ4, RQ5 and RQS.L8B1 completed

  • Long Run including the following circuits launched around 19:30 and ongoing.

RCBYHS4.L8B1 RCBYVS4.L8B2 RCBYVS4.L8B1 RCBYV4.L8B1 RCBYH4.L8B2 RCBCH5.L8B1 RCBCV5.L8B2

RCBV23.L8B2 RCBH23.L8B1 RCBH25.L8B1 RCBV27.L8B2 RCBH27.L8B1 RCBV29.L8B2 RCBH31.L8B1 RCBV32.L8B1 RCBH17.L8B1 RCBV17.L8B2 RCBH16.L8B2 RCBV14.L8B1 RCBH13.L8B1 RCBV13.L8B2 RCBH12.L8B2 RCBV11.L8B2 RCBV24.L8B1 RCBH24.L8B2 RCBV25.L8B2 RCBV28.L8B1 RCBH28.L8B2 RCBH29.L8B1 RCBV31.L8B2 RCBH32.L8B2 RCBH18.L8B2 RCBV18.L8B1 RCBV16.L8B1 RCBH15.L8B1 RCBV15.L8B2 RCBH14.L8B2 RCBV12.L8B1 RCBH11.L8B1

RQD.A78 RQF.A78 RB.A78 RD2.L8 RQ4.L8B2 RQ4.L8B1 RQ5.L8B1 RQ5.L8B2

RQS.L8B1

Test Plan for Wednesday 04.07.2007

See attached file

Maria Paz Casas, Antonio Vergara and Blanca Perea for Hardware Commissioning- 03 Jul 2007

Edit, attach file or add comment


Created by Main.AntonioVergara on 2007-07-03 - 16:39

ElQA Sector 7-8 - Last Week of Powering Tests

Microsoft Excel Spreadsheet ElQAS78.xls:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x07x03x16x21/ElQAS78.xls

Comments:

Edit, attach file or add comment


Created by Main.AntonioVergara on 2007-07-03 - 08:52

Powering Tests - Sector 7-8 Morning Plan

Microsoft Excel Spreadsheet PlanningW27_030707.xls:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x07x03x08x51/PlanningW27_030707.xls

Comments:

Edit, attach file or add comment


Created by Main.AntonioVergara on 2007-07-02 - 22:12

Powering Tests Sector 7-8 - Report 02.07.2007

Microsoft Excel Spreadsheet PlanningW27_020707.xls:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x07x02x21x18/PlanningW27_020707.xls

Test Report from Monday 02.07.2007

* RQF.A78, RQF.A78 and RB.A78: Stability tests completed.

* RCBCH5.L8B1 stability test completed.

* 60A Closed Orbit Correctors: PNO completed for all the circuits released by ElQA. Tests OK except for two natural quenches in the circuits: RCBH28.L8B2 and RCBH25.L8B1

* 600A Correctors:

  • PCC and PIC2 completed for all the circuits available in DFBAO except RQTD.A78B1 and RQTD.A78B2
  • PCS completed successfully for RQTF.A78B1 and RQTF.A78B2
  • RCS.A78B2 - PCS.2 (50A) PCS.3 (100A) - Test OK
  • RCS.A78B1 - PCS.2 failed due to inversed polarity

Test Plan for Tuesday 03.07.2007

See file attached

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-07-02 - 12:11

Powering Test Sector 78 - 2nd July 2007

Microsoft Excel Spreadsheet PlanningW27_020707_.xls:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x07x02x12x09/PlanningW27_020707_.xls

Comments:

Edit, attach file or add comment


Created by Main.MariaPazCasasLino on 2007-07-02 - 11:06

Short Circuit Test in UA23. Minutes 02-07-2007

PDF RAT_20070702.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint2x2007x07x02x11x05/RAT_20070702.pdf

Comments:

Edit, attach file or add comment


Created by Main.JacekSzkutnik on 2007-07-02 - 11:04

Cooldown Sector 78

We are now in the last week before the warm-up of sector 78 for the repair and connection of the inner triplet magnets and some consolidation work which is expected to last six weeks.

During the past week, the commissioning of sector 78 has continued on several fronts: the main magnet circuits, the 600 A corrector circuits, the 60A dipole correctors as well as the matching section magnets. This activity was made possible thanks to the very stable operation of the cryogenic system and of the utilities almost throughout the week. Unfortunately starting from Thursday morning two independent faults (return heater, 24V power supply) in the cryogenic system compromised the powering tests at 1.9K for the rest of the week.

After the two quadrupole circuits, which were both powered to 6.5 kA, the dipole circuit was successfully taken to 2 kA on Wednesday (June 27th) morning. The first cycle is shown below. This cycle ends with a discharge of the circuit in the dump resistor.

The in depth analysis of the data collected during these runs is ongoing to validate the correct operation of all the equipment.

The commissioning of the 600A circuit could finally start after the careful tuning of the quench detection system and the successful interlock test with the power converters connected to the magnets. The quench detection system is set to run with very low detection thresholds. The tests on these circuits are aimed at validating the intermagnet splices some of which could have been executed not to the required standards. These tests will continue this week.

Although not at full thrust, the commissioning of the 60 A orbit corrector circuits is progressing. In addition to all this, detailed studies on many circuits were performed to better identify the powering parameters of these circuits.

Studies to better understand the coupling between the Q4 and the D2 magnets did not succeed to reproduce the phenomenon which was initially observed despite all the efforts.

The preparation for the cool down of sector 45 is well advanced: the electrical quality assurance identified and localized a short in the dipole circuit which was very quickly repaired. A mechanical non conformity was also detected during the flushing: it concerned the swap of two lines in the interconnect between the low and high current module of the DFBAH. This was also quickly repaired and the conditioning of the lines continued over the weekend.

The puppets of the Joule-Thomson CV910 were also replaced and the cooldown is expected to start on Wednesday July 4th.

PNG Cooldown_78_Jul02_1.PNG:
Cooldown_78_Jul02_1.PNG

PNG Cooldown_78_Jul02_2.PNG:
Cooldown_78_Jul02_2.PNG

Comments:

Edit, attach file or add comment

Edit | Attach | Watch | Print version | History: r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r1 - 2007-07-02 - JacekSzkutnik
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    HCC All webs login